OpenStudy (anonymous):

What is the sum of the arithmetic sequence 308, 304, 300, …, if there are 21 terms?

OpenStudy (anonymous):

sum= 21(2*308+(21-1)(-4))/2 sum=21(616-80)/2 sum=5628